OpenStudy (anonymous):

What is the solution of the equation? (x-2)^3/4 = 8

OpenStudy (anonymous):

is that \[\frac{ (x - 2)^{3} }{ 4 } = 8\] or \[(x - 2)^{3/4} = 8\]

OpenStudy (anonymous):

the second one

OpenStudy (anonymous):

all right so the left hand side is raised to the 3/4 power. is there a way to invert that? here is an example: \[y^{5/7} = 10 \rightarrow (y^{5/7})^{7/5} = 10^{7/5} \rightarrow y = 10^{7/5}\]

OpenStudy (anonymous):

i honestly have no clue what to do..

OpenStudy (anonymous):

raise both sides to the 4/3 power. that will cancel the exponents on the left side

OpenStudy (anonymous):

so it would look like this: \[(x-2) = 8_{4}^{3}\]

OpenStudy (anonymous):

we are raising both sides to the 4/3 power so . . .

OpenStudy (anonymous):

\[(x - 2) = 8^{4/3}\]

OpenStudy (anonymous):

sorry i got the exponent flipped around... So now what would be the next step?

OpenStudy (anonymous):

We can focus on analyzing 8^(4/3). another way of writing 8 to the 4/3 is saying the 3rd root of 8^4 \[8^{4/3} = \sqrt[3]{8^{4}}\]

OpenStudy (anonymous):

and without using a calculator .. we could also take note that 8 = 2^3 and rewrite the expression as: \[\sqrt[3]{8^{4}} = \sqrt[3]{(2^{3})^{4}} = \sqrt[3]{2^{3} * 2^{3} * 2^{3} * 2^{3} }\]

OpenStudy (anonymous):

taking the cubed root of each factor here: \[\sqrt[3]{2^{3} * 2^{3} * 2^{3} * 2^{3}} = 2 * 2 * 2 * 2 = 16\]

OpenStudy (anonymous):

so, when it is all said and done \[8^{4/3} = 16\] and our equation that was (x - 2) = 8^(4/3) is now simplified to (x - 2) = 16

OpenStudy (anonymous):

and then you just add 2 to both sides for the final answer?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

thank you i appreciate it. i'm gonna write this in my notes so i can remember all that lol

OpenStudy (anonymous):

i mean if you could use a calculator it would be a lot easier, but i wasnt sure so i took the long route